find the number of millimoles of solute in:



a. 2.00 L of 2.76 x 10-3 M KMnO4


b. 250 mL of a solution that contains 4.20 ppm of CuSO4

1
Expert's answer
2022-02-16T07:12:05-0500

a) CM = n / V

n = CM x V = 2.76 x 10-3 x 2.00 = 5.52 x 10-3 mol = 5.52 mmol


b) ppm = mg/L

m (CuSO4) = 4.20 x 0.25 = 1.05 mg = 0.00105 g

n = m / M

M (CuSO4) = 159.6 g/mol

n (CuSO4) = 0.00105 / 159.6 = 0.00000658 mol = 0.00000000658 mmol